Is a triangle with sides9 9 units, 16units and 25 units a right triangle
Found 3 solutions by mananth, richard1234, ikleyn:
Answer by mananth(16949)

If it is a right triangle then
by pythagoras theorem
25^2= 16^2+9^2
625 = 256+81
625 is not equal to 337
so it is not a right triangle

Answer by richard1234(7193)

No. A triangle with sides 9, 16, 25 is not even a triangle, since 9 + 16 equals 25, which does not satisfy the triangle inequality.

Answer by ikleyn(53996)

.
Is a triangle with sides 9 units, 16 units and 25 units a right triangle ?
~~~~~~~~~~~~~~~~~~~~~~~~~~


Good question.   It is a huge trap,  instead.

        Notice that   9 + 16 = 25,

so the triangle inequality is violated,  which means that a triangle
with the side lengths as described  DOES  NOT  EXIST  and  CAN  NOT  EXIST.


                    Ha - ha - ha.